OpenStudy (anonymous):

What three consecutive odd integers add up to -27?

OpenStudy (stamp):

Consecutive integers are separated by + 2. 2, 4, 6 are all 2 away. 1, 3, 5 are all 2 away integer a = x integer b = ( x ) + 2 = x + 2 integer c = ( x + 2 ) + 2 = x + 4 integer( a + b + c ) = - 27 x + ( x + 2 ) + ( x + 4 ) = - 27 Simplify. 3x + 6 = - 27 3x = - 33 x = -11 a = - 11 b = - 9 c = - 7 Also, -11 + -9 + -7 = - 27

OpenStudy (anonymous):

*consecutive odd/even integers* are separated by 2.